# Large-Area Fabrication-aware Computational Diffractive Optics

### v1

@@ Abstract.
As such, we aim at lifting the fundamental technical barriers to the practical use of learned diffractive optical systems. To this end, we propose a fabrication-aware design pipeline for diffractive optics fabricated by direct-write grayscale lithography followed by replication with nano-imprinting, which is directly suited for inexpensive mass-production of large area designs. We propose a super-resolved neural lithography model that can accurately predict the 3D geometry generated by the fabrication process. This model can be seamlessly integrated into existing differentiable optics frameworks, enabling fabrication-aware, end-to-end optimization of computational optical systems across a wide range of applications.
To tackle the computational challenges of such large-scale inverse designs at high resolution, especially the high memory consumption, we also devise tensor-parallel compute framework centered on distributing large-scale FFT computation across many GPUs.
Using this combination of methods, we demonstrate large scale diffractive optics designs up to 32.16 mm 21.44 mm, simulated on grids of up to 128,640 by 85,760 feature points. We find adequate agreement between simulation and fabricated prototypes for applications such as holography and PSF engineering. We also achieve high image quality from an imaging system comprised only of a single diffractive optical element, with images processed only by a one-step inverse filter utilizing the simulation PSF. We believe our findings lift the fabrication limitations for real-world applications of diffractive optics and differentiable optical design.
@@ 1. Introduction
To understand the reason for the discrepancy between simulated designs and fabricated prototypes, we need to consider a specific fabrication process in some detail. Here we explain grayscale lithography with direct laser writing (see Fig. 1, left); other lithographic method such a two-photon polymerization (Wang et al., 2023) differ in the details but require similar high-level considerations. In direct laser writing, a laser beam scans across a wafer covered in thin layer photoresist in a 2D grid. Modulation of the beam intensity is used to create a spatially varying exposure map (A). The first notable effect in this process is an optical blur of the specified design, which can be modeled as the convolution of the design pattern with the point spread function of the laser beam and can, for example, result in corner rounding for small rectangular features. Next, the local exposure level induces a local chemical change in the photoresist, which can be thought of as a nonlinear local transfer function. The next step (B) is the development of the exposed resin, in which the developer decomposes the photoresist, where the rate of decomposition depends on the local exposure received by the resist. This is a complicated 3D chemical interaction that finally reveals a height field structure. Nanoimprinting can be used to transfer this shape into a UV resin (C) to form the final DOE (D). Further shape distortions are possible in this step, for example, due to volume shrinkage of the resin during curing. For mass production, only the final imprinting step needs to be repeated for each copy.
@@
A related source of error is that most inverse design pipelines for diffractive optics work on drastically undersampled grids out of computational necessity. Most early DOE design works use simulation grids where one grid point equals one DOE feature. More recently, simulations have started using a finer grid (e.g. (Yang et al., 2024b)). However, according to the Nyquist limit any sample grid can only represent sinusoidal components up to twice the grid pitch, whereas many holgraphy and DOE designs rely on sharp step edges between neighboring features. We experimentally demonstrate this problem in Section 3.2 and Figure 3 by performing simulations on the original feature grid resolution and on an super-resolved grid with nearest neighbor upsampling to force box-shaped features, highlighting drastic simulation differences even without accounting for fabrication limitations.
Several strategies for dealing with the large simulation gap have been introduced and are in common use: commercial product design often relies on a tedious iterative process between design, fabrication, and measurement (Jang et al., 2020). In the holographic display community, camera-in-the-loop systems have recently become popular (Peng et al., 2020), whereas in end-to-end designed computational imaging systems, it is common practice to fine-tune the computational module with measurements from the as-fabricated prototype (e.g. (Peng et al., 2019; Shi et al., 2024a; Chakravarthula et al., 2023) etc.). However, closing the loop in this fashion is expensive and may not be practical beyond lab prototypes.
@@
The last application confirms the ability to not only eliminate fine-tuning of the image restoration module based on measured characterization of the prototype, but also demonstrates that, with a design process that includes an accurate model of the manufacturing process, image restoration does not necessarily require heavy deep neural networks but can utilize more lightweight architectures that are compatible with the computational resources of edge devices. We believe that this, by itself, is a major step forward in improving the practicality of diffractive end-to-end design beyond lab prototypes.
@@ 2.1. Diffractive Optical Elements
In 1948, Dennis Gabor introduced the optical holography, the first physical realization of DOEs through interference (Gabor, 1948). Since then, various forms of DOEs have been proposed, such as computer-generated holograms (Brown and Lohmann, 1966), binary gratings (Collischon et al., 1994) and kinoform lens (Lesem et al., 1969), to name a few—see (Zhang et al., 2023a) for a comprehensive review of long history of DOEs. However, due to intrinsic difficulty in designing and manufacturing DOEs, their uses remain largely limited to well-controlled laboratory settings. Over the last decade, advances in computational imaging and the increasing capabilities of nano-fabrication technology have allowed researchers to revisit these conventional DOEs, as powerful optical encoders in hardware-software co-designed computational systems (Peng et al., 2015, 2016; Heide et al., 2016; Asif et al., 2017; Boominathan et al., 2016; Sitzmann et al., 2018). These methods have been successful in a wide array of applications, including full-spectral color imaging (Peng et al., 2016; Heide et al., 2016), extended depth-of-field imaging (Tan et al., 2021), compressive lensless imaging (Asif et al., 2017; Boominathan et al., 2020), single-shot hyperspectral (Shi et al., 2024b) and depth imaging (Baek et al., 2021; Shi et al., 2024a), high-dynamic-range imaging (Metzler et al., 2020; Sun et al., 2020), seeing through obstructions (Shi et al., 2022), ultra-wide-angle holographic display (Tseng et al., 2023), and computer vision tasks (Wei et al., 2024). Although successful, two fundamental technical obstacles remain: 1) the gap from design to manufacturing and 2) the challenge to design large-area devices at high fidelity, which are essential to compete with the widely used refractive optics.
@@ 2.3. Parallel Computing and Distributed Training in Deep Learning
Our work also takes inspiration from recent advances in parallel computing and distributed training infrastructure (Radford et al., 2019; Brown et al., 2020; Narayanan et al., 2021). The most common parallelization strategy is data parallelism (Hillis and Steele Jr, 1986), which distributes the data across different computing nodes and operate on the data in parallel. However, this technique has a fundamental limitation in the model size it can tackle—the model must fit entirely on one worker. With the increasing size and complexity, NNs have approached the memory capacity of modern hardware accelerators. To overcome this bottleneck, model parallelism (Shoeybi et al., 2020) has been proposed for training billion-parameter-scale LLMs, which includes pipeline parallelism (Huang et al., 2019) and more general tensor parallelism (Narayanan et al., 2021). Pipeline parallelism splits the NN pipeline into multiple stages across devices at the expense of the bubble overhead; Tensor parallelism is specialized for distributing specific atomic operators, such as the general matrix multiplication widely used in the Transformer block (Vaswani et al., 2017). For computational optics, data parallelism supports scaling of wavelengths, incident angles, propagation distances, but cannot realize designs of large-area/scale DOEs discretized with billion-pixel-scale huge arrays. As such, we leverage the tensor parallelism and devise the specialized computing routines for the large-area wave-optics model. Our implementation follows the generalized single program multiple data (GSPMD) programming model (Xu et al., 2021; Shazeer et al., 2018), and thereby flexibly supports any tensor-dimensions to be split across any dimensions of a multi-dimensional mesh of processors—realizing arbitrary hybrid data and model (tensor) parallelism without painful code rewriting.
@@ 3.2. Super-resolved Neural Lithography Model
A 3.6-mm2 DOE with a feature size of 2 m is optimized to generate a 2-D hologram image under coherent (520.6 nm wavelength) laser illumination. We use the Adam optimizer (Kingma, 2014) to solve (4), where the penalty is defined as a combination of a scale-invariant mean square error loss and an energy regularization 111Detailed definitions of the loss functions are given in the Supplementary Material as
where is empirically set as . The DOE height map is randomly initialized, and the learning rate is initially set as , which follows a cosine schedule that decays to zero in 3000 iterations. Such an optimization leads to almost perfect 2-D hologram reconstruction (34.9 dB in PSNR) evaluated under this conventional setting (see the left-top image in Fig. 3). However, as explained in the introduction, the optical system simulation at -spacing grid violates the Nyquist sampling theorem (Smith, 1999), which inevitably results in aliasing artifacts degrading performance in reality. The zero-order interpolation, i.e., nearest upsampling is typically employed (Tseng et al., 2023; Kuo et al., 2023) to alleviate this issue, but this assumes perfect flat feature structure can be reliably manufactured. Other smooth/low-pass interpolation kernels such as the Lanczos-3 kernel (Getreuer, 2011) remain coarse approximations of the de-facto ”fabrication kernel”. Fig. 3 summarizes the results of DOEs optimized and evaluated at combinations of different settings, suggesting that 1) hologram can only be well reconstructed when train and evaluate under the same setting; 2) interpolation kernels play a critical role in high-frequency hologram feature construction, where a mismatched kernel can result in drastic quality decline. As such, this simple yet informative experiment emphasizes the importance of the ”fabrication kernel” and thus motivates the need of fabrication-aware optimization.
@@
This fabrication pipeline offers several advantages over the conventional multi-level binary-mask-based photolithography process, which was widely used in DOE manufacturing in literature (Khonina et al., 2024). Unlike the binary-mask approach, which requires multiple photolithography steps with separate masks for discrete height levels, grayscale lithography uses a single laser exposure to create continuous 3D relief profiles, significantly reducing process complexity and eliminating alignment errors. The simplified process allows for faster prototyping and design iterations, shortening the prototyping cycle from one week to a few hours for a centermeter-scale DOE design, according to our experience in manufacturing.
@@
After contrast curve calibration, we design and fabricate another set of calibration patterns (spatially arranged in a single wafer) to construct the dataset for training and evaluation of the neural lithography model. Similar to prior work (Zheng et al., 2023), we randomly generate a set of 2-D patterns following a uniform distribution. A low-pass filter is then applied in the Fourier domain to limit the high-frequency components. patterns in total are created with incrementally decreasing maximum cutoff frequency, each of which features 2 area discretized at -spacing grid (c.f., Fig. 5). Once fabricated, the resulting sample is then measured using an atomic force microscope (AFM) (Giessibl, 2003) to obtain precise 3-D profiles of the fabricated patterns, which can be used as ground truth for training and evaluating the neural lithography model. For each pattern we scan an area of , with 256256 sampling points, resulting in approximately sampling resolution. To cope with the AFM imaging artifacts inherent in the measurement process (Ricci and Braga, 2004), we measure each pattern twice in two orthogonal scanning directions. The resulting raw measurements are then preprocessed, fused and registered to obtain final 3-D profiles with high fidelity.
@@ 4.1. Large-scale Wave Propagation via D2FFT
Multidimensional FFT can be efficiently computed by a sequence of lower-dimensional FFTs. For instance, a 2-D FFT can be performed by first applying a 1-D FFT along the rows, followed by a 1-D FFT along the columns. Such a row-column decomposition naturally leads to a distributed-memory FFT implementation with an additional communication step. To execute a large-scale 2-D FFT, we first partition the given 2-D array into multiple sub-array chunks along the columns (Y-axis), each of which is assigned and allocated to a different (GPU) device. Under this arrangement, the row (X-axis) information is completely preserved in each device, and hence, a batched 1-D FFT along rows (X-axis) can be performed independently on each device. Then, a MPI-style primitive for collective communication, all-to-all communication (Snir, 1998; Alabed et al., 2025), is invoked to exchange unique chunks of the distributed array between participating devices. Mathematically, this is analogous to a matrix transpose operation, such that the resulting array is now partitioned along the rows (X-axis), which is followed by a batched 1-D FFT along the Y-axis on each device to complete the 2-D FFT. Note the spatial sharding axis of the distributed array swaps after performing this distributed-memory 2-D FFT. The whole process is illustrated in Fig. 6 (top-left). We implement this operation in Jax (Bradbury et al., 2018) framework in a differentiable way, where the backward pass is realized by another distributed-memory 2-D FFT since the discrete Fourier transform can be represented by a symmetric matrix. Differentiable distributed-memory inverse FFT can be implemented similarly.
@@ 4.2. Tensor-parallel Convolutional Neural Network
[ frame=lines, framesep=2mm, baselinestretch=1.2, fontsize=]python ””” A multi-dimensional processor mesh is created with 2x2 devices where first and second axes are named wavelength (”wvl”) and tensor parallel (”tp”) respectively. ””” mesh = MeshShardingHelper([2, 2], [’wvl’, ’tp’]) @partial( mesh.sjit, args_sharding_constraint=( PartitionSpec(’wvl’, ’tp’, None), # for ”field” PartitionSpec(’wvl’, None, ’tp’), # for ”H” ), out_shardings=PartitionSpec(’wvl’, ’tp’, None), # for ”E_prop” ) def propagate(field, H): U = fft_func(field) # Perform 2D FFT (implemented elsewhere) E_k_prop = U * H E_prop = ifft_func(E_k_prop) # Perform 2D inverse FFT return E_prop

### v2

@@ Abstract.
We aim at lifting the fundamental technical barriers to the practical use of learned diffractive optical systems. To this end, we propose a fabrication-aware design pipeline for diffractive optics fabricated by direct-write grayscale lithography followed by replication with nano-imprinting, which is directly suited for inexpensive mass-production of large area designs. We propose a super-resolved neural lithography model that can accurately predict the 3D geometry generated by the fabrication process. This model can be seamlessly integrated into existing differentiable optics frameworks, enabling fabrication-aware, end-to-end optimization of computational optical systems. To tackle the computational challenges, we also devise tensor-parallel compute framework centered on distributing large-scale FFT computation across many GPUs.
As such, we demonstrate large scale diffractive optics designs up to 32.16 mm 21.44 mm, simulated on grids of up to 128,640 by 85,760 feature points. We find adequate agreement between simulation and fabricated prototypes for applications such as holography and PSF engineering. We also achieve high image quality from an imaging system comprised only of a single diffractive optical element, with images processed only by a one-step inverse filter utilizing the simulation PSF. We believe our findings lift the fabrication limitations for real-world applications of diffractive optics and differentiable optical design.
@@ 1. Introduction
To understand the reason for the discrepancy between simulated designs and fabricated prototypes, we need to consider a specific fabrication process in some detail. Here we explain grayscale lithography with direct laser writing (see Fig. 1, left) (Grushina, 2019); other lithographic method such a two-photon polymerization (Wang et al., 2023) differ in the details but require similar high-level considerations. In direct laser writing, a laser beam scans across a wafer covered in thin layer photoresist in a 2D grid. Modulation of the beam intensity is used to create a spatially varying exposure map (A). The first notable effect in this process is an optical blur of the specified design, which can be modeled as the convolution of the design pattern with the point spread function of the laser beam and can, for example, result in corner rounding for small rectangular features. Next, the local exposure level induces a local chemical change in the photoresist, which can be thought of as a nonlinear local transfer function. The next step (B) is the development of the exposed resin, in which the developer decomposes the photoresist, where the rate of decomposition depends on the local exposure received by the resist. This is a complicated 3D chemical interaction that finally reveals a height field structure. Nanoimprinting can be used to transfer this shape into a UV resin (C) to form the final DOE (D). Further shape distortions are possible in this step, for example, due to volume shrinkage of the resin during curing. For mass production, only the final imprinting step needs to be repeated for each copy.
@@
A related source of error is that most inverse design pipelines for diffractive optics work on drastically undersampled grids out of computational necessity. Most early DOE design works use simulation grids where one grid point equals one DOE feature. More recently, simulations have started using a finer grid (e.g., (Yang et al., 2024b)). However, according to the Nyquist limit any sample grid can only represent sinusoidal components up to twice the grid pitch, whereas many holography and DOE designs rely on sharp step edges between neighboring features. We experimentally demonstrate this problem in Section 3.2 and Figure 3 by performing simulations on the original feature grid resolution and on an super-resolved grid with nearest neighbor upsampling to force box-shaped features, highlighting drastic simulation differences even without accounting for fabrication limitations.
Several strategies for dealing with the large simulation gap have been introduced and are in common use: commercial product design often relies on a tedious iterative process between design, fabrication, and measurement (Jang et al., 2020). In the holographic display community, camera-in-the-loop systems have recently become popular (Peng et al., 2020; Kavaklı et al., 2021; Choi et al., 2022), whereas in end-to-end designed computational imaging systems, it is common practice to fine-tune the computational module with measurements from the as-fabricated prototype (e.g., (Peng et al., 2019; Shi et al., 2024a; Chakravarthula et al., 2023) etc.). However, closing the loop in this fashion is expensive and may not be practical beyond lab prototypes.
@@
The last application confirms the ability to not only eliminate fine-tuning of the image restoration module based on measured characterization of the prototype, but also demonstrates that, with a design process that includes an accurate model of the manufacturing process, image restoration does not necessarily require heavy deep neural networks but can utilize more lightweight architectures that are compatible with the computational resources of edge devices. We believe that this, by itself, is a major step forward in improving the practicality of diffractive end-to-end design beyond lab prototypes. Our code is publicly available at https://github.com/Vandermode/LAFA
@@ 2.1. Diffractive Optical Elements
In 1948, Dennis Gabor introduced the optical holography, the first physical realization of DOEs through interference (Gabor, 1948). Since then, various forms of DOEs have been proposed, such as computer-generated holograms (Brown and Lohmann, 1966), binary gratings (Collischon et al., 1994) and kinoform lens (Lesem et al., 1969), to name a few—see (Zhang et al., 2023a) for a comprehensive review of long history of DOEs. However, due to intrinsic difficulty in designing and manufacturing DOEs, their uses remain largely limited to well-controlled laboratory settings. Over the last decade, advances in computational imaging and the increasing capabilities of nano-fabrication technology have allowed researchers to revisit these conventional DOEs, as powerful optical encoders in hardware-software co-designed computational systems (Peng et al., 2015, 2016; Heide et al., 2016; Asif et al., 2017; Boominathan et al., 2016; Sitzmann et al., 2018; Wu et al., 2019). These methods have been successful in a wide array of applications, including full-spectral color imaging (Peng et al., 2016; Heide et al., 2016), extended depth-of-field imaging (Tan et al., 2021), compressive lensless imaging (Asif et al., 2017; Boominathan et al., 2020), single-shot hyperspectral (Shi et al., 2024b) and depth imaging (Baek et al., 2021; Shi et al., 2024a), high-dynamic-range imaging (Metzler et al., 2020; Sun et al., 2020), seeing through obstructions (Shi et al., 2022), ultra-wide-angle holographic display (Tseng et al., 2024), and computer vision tasks (Wei et al., 2024). Although successful, two fundamental technical obstacles remain: 1) the gap from design to manufacturing and 2) the challenge to design large-area devices at high fidelity, which are essential to compete with the widely used refractive optics.
@@ 2.3. Parallel Computing and Distributed Training in Deep Learning
Our work also takes inspiration from recent advances in parallel computing and distributed training infrastructure (Radford et al., 2019; Brown et al., 2020; Narayanan et al., 2021). The most common parallelization strategy is data parallelism (Hillis and Steele Jr, 1986), which distributes the data across different computing nodes and operate on the data in parallel. However, this technique has a fundamental limitation in the model size it can tackle—the model must fit entirely on one worker. With the increasing size and complexity, NNs have approached the memory capacity of modern hardware accelerators. To overcome this bottleneck, model parallelism (Shoeybi et al., 2020) has been proposed for training billion-parameter-scale LLMs, which includes pipeline parallelism (Huang et al., 2019) and more general tensor parallelism (Narayanan et al., 2021). Pipeline parallelism splits the NN pipeline into multiple stages across devices at the expense of the bubble overhead; Tensor parallelism is specialized for distributing specific atomic operators, such as the general matrix multiplication widely used in the Transformer block (Vaswani et al., 2017). For computational optics, data parallelism supports scaling of wavelengths, incident angles, propagation distances, but cannot realize designs of large-area/scale DOEs discretized with billion-pixel-scale huge arrays (Sun et al., 2025). As such, we leverage the tensor parallelism and devise the specialized computing routines for the large-area wave-optics model. This includes differentiable distributed-memory (D2) FFT for free-space wave propagation as well as spatial-partitioning convolutional neural networks as the neural lithography model. Although the basic idea of distributed-memory FFT can be traced back to decades ago in the high-performance computing community (Gupta and Kumar, 2002) with implementations in CPU (Frigo and Johnson, 2005; Pippig, 2013), GPU (Gholami et al., 2016; Ayala et al., 2022) or TPU (Lu et al., 2021) clusters based on Message Passing Interface standard (MPI) (Snir, 1998), none of these methods have been designed for differentiable wave optics simulation. Our implementation follows the generalized single program multiple data (GSPMD) programming model (Xu et al., 2021; Shazeer et al., 2018), and thereby flexibly supports any tensor-dimensions to be split across any dimensions of a multi-dimensional mesh of processors—realizing arbitrary hybrid data and model (tensor) parallelism without painful code rewriting.
@@ 3.2. Super-resolved Neural Lithography Model
A 3.63.6 mm2 DOE with a feature size of 2 m is optimized to generate a 2-D hologram image under coherent (520.6 nm wavelength) laser illumination. We use the Adam optimizer (Kingma, 2014) to solve (4), where the penalty is defined as a combination of a scale-invariant mean square error loss and an energy regularization 111Detailed definitions of the loss functions are given in the Supplementary Material as
where is empirically set as . The DOE height map is randomly initialized, and the learning rate is initially set as , which follows a cosine schedule that decays to zero in 3000 iterations. Such an optimization leads to almost perfect 2-D hologram reconstruction (34.9 dB in PSNR) evaluated under this conventional setting (see the left-top image in Fig. 3). However, as explained in the introduction, the optical system simulation at -spacing grid violates the Nyquist sampling theorem (Smith, 1999), which inevitably results in aliasing artifacts degrading performance in reality. The zero-order interpolation, i.e., nearest upsampling is typically employed (Kuo et al., 2023; Tseng et al., 2024) to alleviate this issue, but this assumes perfect flat feature structure can be reliably manufactured. Other smooth/low-pass interpolation kernels such as the Lanczos-3 kernel (Getreuer, 2011) remain coarse approximations of the de-facto ”fabrication kernel”. Fig. 3 summarizes the results of DOEs optimized and evaluated at combinations of different settings, suggesting that 1) hologram can only be well reconstructed when train and evaluate under the same setting; 2) interpolation kernels play a critical role in high-frequency hologram feature construction, where a mismatched kernel can result in drastic quality decline. As such, this simple yet informative experiment emphasizes the importance of the ”fabrication kernel” and thus motivates the need of fabrication-aware optimization.
@@
This fabrication pipeline offers several advantages over the conventional multi-level binary-mask-based photolithography process, which was widely used in DOE manufacturing in literature (Khonina et al., 2024). Unlike the binary-mask approach, which requires multiple photolithography steps with separate masks for discrete height levels, grayscale lithography uses a single laser exposure to create continuous 3-D relief profiles, significantly reducing process complexity and eliminating alignment errors. The simplified process allows for faster prototyping and design iterations, shortening the prototyping cycle from one week to a few hours for a centimeter-scale DOE design, according to our experience in manufacturing.
@@
After contrast curve calibration, we design and fabricate another set of calibration patterns (spatially arranged in a single wafer) to construct the dataset for training and evaluation of the neural lithography model. Similar to prior work (Zheng et al., 2023), we randomly generate a set of 2-D patterns following a uniform distribution. A low-pass filter is then applied in the Fourier domain to limit the high-frequency components. patterns in total are created with incrementally decreasing maximum cutoff frequency, each of which features 2 area discretized at -spacing grid (c.f., Fig. 5). Once fabricated, the resulting sample is then measured using an atomic force microscope (AFM) (Giessibl, 2003) to obtain precise 3-D profiles of the fabricated patterns, which can be used as ground truth for training and evaluating the neural lithography model. For each pattern we scan an area of , with 256256 sampling points, resulting in approximately sampling resolution. To cope with the AFM imaging artifacts inherent in the measurement process (Ricci and Braga, 2004), we measure each pattern twice in two orthogonal scanning directions. The resulting raw measurements are then preprocessed, registered and fused to obtain final 3-D profiles with high fidelity.
@@ 4.1. Large-scale Wave Propagation via D2FFT
Multidimensional FFT can be efficiently computed by a sequence of lower-dimensional FFTs. For instance, a 2-D FFT can be performed by first applying a 1-D FFT along the rows, followed by a 1-D FFT along the columns. Such a row-column decomposition naturally leads to a distributed-memory FFT implementation with an additional communication step. To execute a large-scale 2-D FFT, we first partition the given 2-D array into multiple sub-array chunks along the columns (Y-axis), each of which is assigned and allocated to a different (GPU) device. Under this arrangement, the row (X-axis) information is completely preserved in each device, and hence, a batched 1-D FFT along rows (X-axis) can be performed independently on each device. Then, a MPI-style primitive for collective communication, all-to-all communication (Snir, 1998; Doi and Negishi, 2010; Alabed et al., 2025), is invoked to exchange unique chunks of the distributed array between participating devices. Mathematically, this is analogous to a matrix transpose operation, such that the resulting array is now partitioned along the rows (X-axis), which is followed by a batched 1-D FFT along the Y-axis on each device to complete the 2-D FFT. Note the spatial sharding axis of the distributed array swaps after performing this distributed-memory 2-D FFT. The whole process is illustrated in Fig. 6 (top-left). We implement this operation in Jax (Bradbury et al., 2018) framework in a differentiable way, where the backward pass is realized by another distributed-memory 2-D FFT since the discrete Fourier transform can be represented by a symmetric matrix. Differentiable distributed-memory inverse FFT can be implemented similarly.